OpenStudy (anonymous):

Use implicit differentiation to find an equation of the tangent line to the curve at the given point. 2(x^2+y^2)^2=25(x^2−y^2)at point (3,1)

OpenStudy (anonymous):

what did you get for \(\Large{dy\over dx}\)

OpenStudy (anonymous):

well i wasn't sure if i had to do the product rule but i got 8x^3 +4x^2(2y)dy/dx + 8xy^2+8y^3 dy/dx= 50x-50y dy/dx

OpenStudy (anonymous):

good\[ 4(x^2+y^2)\left(2x+2y{dy\over dx}\right)=25\left(2x-2y{dy\over dx}\right) \] rearrange the terms so that you get something like this: \[ {dy\over dx}={\rm something} \]

OpenStudy (anonymous):

would i have to use the product rule or just distribute

OpenStudy (anonymous):

just distribute.

OpenStudy (anonymous):

since there are no two functions multiplying, you do not have to use product rule.

OpenStudy (anonymous):

im not sure if i did the math correctly but i got dy/dx= 21/608

OpenStudy (anonymous):

Who deleted my post lol

OpenStudy (anonymous):

\[ {dy\over dx}=\frac{25x-4x(x^2+y^2)}{25y+4y(x^2+y^2)} \] and plug in x=3, y=1

OpenStudy (anonymous):

is that what you did?

OpenStudy (anonymous):

yeah and for that i got 7/ 13 is that correct

OpenStudy (anonymous):

positive?

OpenStudy (anonymous):

umm no i got -7/13

OpenStudy (anonymous):

should be -9/13

OpenStudy (anonymous):

maybe i subtracted wrong but yeah you are correct thank you :)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

what did you get for the y-intercept,? what is the equation of tangent line?

OpenStudy (anonymous):

for my final equation i got y=-9/13x+40/13